OpenStudy (anonymous):

Find the general solutions x .t/ to the following differential equations: dx/dt =2 (t/x)( looks like a mixed fraction) pls help me to solve it

OpenStudy (anonymous):

hey uthere

terenzreignz (terenzreignz):

I wonder... \[\Large \frac{dx}{dt}= 2 \frac{t}x\] Like this?

OpenStudy (anonymous):

yes... and i got answer \[+-\sqrt{\frac{ 2 }{ 1} t2+c}\]. is it correct?